ns Here is a bivariate data set. 49.3 155.2 6.7 32.3 47.6 $0.9 44.3 44.8 (0/10) (0/10) [10/10] 6 (0/10) 298 5.7 15.6 -30.2 33.2 -16.8 27.9 45.9 21 22.1 39.2 -3.1 46.4 -113.1 11 (2/10) 212 (2/10) 13 (0/10) 214 (0/10) Q15 (0/10) Q 16 (0/10) Find the correlation coefficient and report it accurate to three decimal places. 17 (0/10) Points possible: 10 Unlimited attempts. Q 19 (0/10) Submit t Version 川1 6
media%2F1a9%2F1a9cdeb5-bd54-4549-b1ec-47
Answer #1

Correlation coefficient:

X Y xy x2 y2
49.3 155.2 7651.36 2430.49 24087.04
6.7 32.3 216.41 44.89 1043.29
47.6 50.9 2422.84 2265.76 2590.81
44.3 44.8 1984.64 1962.49 2007.04
29.8 5.7 169.86 888.04 32.49
15.6 -30.2 -471.12 243.36 912.04
33.2 -16.8 -557.76 1102.24 282.24
27.9 -45.9 -1280.61 778.41 2106.81
21 -22.1 -464.1 441 488.41
39.2 -3.1 -121.52 1536.64 9.61
46.4 -113.1 -5247.84 2152.96 12791.61
total 361 57.7 4302.16 13846.28 46351.39
n 11

r=11*4302.16-(361)(57.7)/\sqrt{[11*13846.28-(361)^{2}][11*46351.39-(57.7)^{2}]}

r=47323.76-20829.7/\sqrt{[152309.08-130321][509865.29-3329.29]}

r=26494.06/\sqrt{[21988.08][506536]}

r=26494.06/105535.5584

r=0.251

Correlation coefficeint:

X Y xy x2 y2
48.4 43.3 2095.72 2342.56 1874.89
35.2 40.7 1432.64 1239.04 1656.49
92.8 253.4 23515.52 8611.84 64211.56
66.7 259.2 17288.64 4448.89 67184.64
61.1 -69.2 -4228.12 3733.21 4788.64
46.8 -68.5 -3205.8 2190.24 4692.25
64.4 23.8 1532.72 4147.36 566.44
62 217.7 13497.4 3844 47393.29
74.9 24.8 1857.52 5610.01 615.04
54.3 -26.1 -1417.23 2948.49 681.21
total 606.6 699.1 52369.01 39115.64 193664.45
n 10

r=10*52369.01-(606.6)(699.1)/\sqrt{[10*39115.64-(606.6^{2}][10*193664.45-(699.1)^{2}]}

r=523690.1-424074.06\sqrt{[391156.4-367963.56][1936644.5-488740.81]}

r=99616.04/\sqrt{[23192.84][1447903.69]}

r=99616.04/183251.1899

r=0.543

what proportion of the variation in y can be explained by the variation in the values of x?

r2=(0.543)2

=0.2948

=0.29

29%proportion of the variation in y can be explained by the variation in the values of x